OpenStudy (anonymous):

Use the function rule: f(x) = 3x - 2 Find the output f(1.5) Will medal and fan :3

OpenStudy (anonymous):

f(x)=3x−2 f(1.5)=3(1.5)−2 f(1.5)=4.5−2 f(1.5)=2.5

OpenStudy (anonymous):

Thank you :) sorry for the delay, I was in an online science class

OpenStudy (anonymous):

lol its all good